OpenStudy (anonymous):

A motorcyclist changes the velocity of his bike from 20.0 meters/second to 35.0 meters/second under a constant acceleration of 4.00 meters/second2. How long does it take the bike to reach the final velocity?

OpenStudy (anonymous):

v=u+at so 35-20/4 = t

OpenStudy (carmuz):

\[V_f = V_0 + at    \rightarrow     35 = 20 + 4t\]